if (x+iy)(p+iq)=(x^2+y^2)i,then, 
a)p=x,q=y
b)p=x^2,q=y^2
c)x=p,y=p
d)x=-q,y=p

Grade:11

2 Answers

Arun
6 years ago
I
Arun
6 years ago
On expansion of LHS
(xp-yq) + (xq+yp)i = (x^2+y^2)i
On comparing real parts
x = yq/p
In comparison of imaginary parts
xq+ yp = x^2+y^2
Put tge value of x = yq/p
y = p
Hence x = q